# A Law of Next-Token Prediction in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have been widely employed across various application domains, yet their black-box nature poses significant challenges to understanding how these models process input data internally to make predictions. In this paper, we introduce a precise and quantitative law that governs the learning of contextualized token embeddings through intermediate layers in pre-trained LLMs for next-token prediction. Our findings reveal that each layer contributes equally to enhancing prediction accuracy, from the lowest to the highest layer—a universal phenomenon observed across a diverse array of open-source LLMs, built on architectures such as Transformer, RWKV, and Mamba. We demonstrate that this law offers new perspectives and insights to inform and guide practices in LLM development and applications, including model scaling, pre-training tasks, and information flow. Overall, our law enables more fine-grained approaches to the design, training, and interpretation of LLMs through scrutinizing their internal data processing mechanisms.
@@
The primary difficulty arises from the hierarchical nature of LLM architectures, such as Transformer [43], RWKV [29], and Mamba [13]. These architectures are composed of multiple layers, each corresponding to simple functions such as linear or quadratic transformations—better known as the attention mechanism—or a nonlinear combination of both. While the input and output can be observed and specified, the internal transformation of data by each layer becomes elusive due to the hierarchical composition. Specifically, in the case of generative pre-trained transformers (GPT) [30], it is unclear how the embeddings of the input text are progressively transformed into features across different layers for next-token prediction, where a token refers to a word or subword. The goal of this paper is to shed light on the inner workings of LLMs as they process token embeddings across all layers. In particular, we aim to identify universal and quantitative patterns that can provide useful principles and insights to refine training processes and enhance interpretability in LLMs.
@@
Fig. 1: The law of equi-learning in large language models. Throughout the paper, the x axis represents the layer index, and the y axis, on a logarithmic scale, represents the PR defined in Eq. 1, unless otherwise specified. The Pearson correlation coefficients, by row first, are -0.997, -0.994, -0.993, -0.988; -0.983, -0.993, -0.992, -0.988; -0.991, -0.989, -0.988, -0.989; -0.988, -0.994, -0.993, -0.992; -0.992, -0.991, -0.991, -0.997. More details can be found in the Supplementary Materials.
Fig. 2: Intermediate-layer contextualized token embeddings for patients</w> (red), cells</w> (blue), and disorder</w> (green) plotted on the plane of the first two principal components.
@@ 1 Main Results
LLMs are nonlinear models that iteratively map a sequence of token embeddings to a new sequence through attention mechanisms and various other operations at each layer. Given a sequence of tokens as input data, a pre-trained LLM transforms these tokens into a sequence of embeddings. The model then uses the last embedding, denoted as $\mathbf{h}_{\text{last}}$, from the last layer to predict the next token in the sequence. Consider $S$ sequences, each containing $T_{s}$ tokens for $1\leq s\leq S$. For $1\leq t\leq T_{s}$, let $x^{s}_{t+1}$ represent the next token that the LLM aims to predict based on the preceding $t$ tokens, $x^{s}_{1}x^{s}_{2}\cdots x^{s}_{t}$, through the last embedding $\mathbf{h}_{t,\text{last}}^{s}$. This process forms a dataset $\mathcal{D}:=\{(\mathbf{h}_{t,\text{last}}^{s},x^{s}_{t+1})\mid 1\leq s\leq S,1\leq t\leq T_{s}-1\}$.
@@
where the sum is over all $x_{\text{next}}=x^{s}_{t+1}$, $\hat{x}_{\text{next}}=\mathbf{w}\cdot\mathbf{h}_{t,\text{last}}^{s}+b$, and $\bar{x}_{\text{next}}$ represents the mean of all $x^{s}_{t+1}$. In statistical terms, this measure is known as the fraction of variance unexplained, equivalent to 1 minus the coefficient of determination. It serves as a canonical metric for evaluating the proportion of variance in the dependent variable that remains unaccounted for by the independent variables. A high PR value indicates limited predictive power of the token embeddings, while a low value suggests strong predictive capability.
To investigate how the predictive power of an LLM with depth $L$ evolves across its layers, we calculate the PR for the next-token prediction task at each intermediate layer. Let $\text{PR}_{l}$ denote this value for the $l$-th layer, where $1\leq l\leq L$. The law of equi-learning (see Fig. 1) states that the dynamics of predictive power across layers follow the relationship
@@
for some decay ratio $10<\rho<$ 1. Since the token embeddings at the input layer (the 0-th layer) are not yet contextualized, they are not included. This implies that $\log\text{PR}_{l+1}-\log\text{PR}_{l}\approx-\log\frac{1}{\rho}$, indicating a roughly constant reduction in the logarithm of the PR value across each layer, hence the name equi-learning. The Pearson correlation coefficients between the logarithm of the PR value and the layer index range from $-0.983$ to $-0.997$ in Fig. 1. For GPT-1, for example, the decay ratio $\rho$ is approximately $0.993$ (as shown in the top-left plot of Fig. 1). In general, $\rho$ depends on various factors, including the model architecture, pre-training data, model depth, feature dimension, and pre-training time.
@@
This law provides what may be the first precise geometric characterization of the learning process for contextualized token embeddings within the intermediate layers of LLMs. Notably, the pre-training objective focuses solely on the last-layer embeddings, aiming to minimize a loss function associated with the last-layer $\text{PR}_{L}$. Surprisingly, this training dynamics inherently ensure that each layer contributes equally, rather than allowing some layers to carry a disproportionate amount of the workload. Related findings have previously been observed in the context of multilayer neural networks [16].
@@
This law is universal, as it emerges across a wide range of open-source LLMs with varying architectures, model sizes, and pre-training data. Additionally, we employ a diverse collection of probing datasets as input to calculate PR and evaluate the law of equi-learning. These datasets include BookCorpus [47], C4 [33], OpenWebText [12], Wikipedia [10], peS2o [37], The Pile [11], Redpajama [6], and OSCAR [38].
When does the law emerge?
@@
To deepen our understanding of the law’s dynamics during training, we investigate the effects of three key factors—training steps, training epochs, and data repetition—on its emergence throughout the process. As illustrated in Fig. 3, the progression of this law as a function of training steps closely resembles the behavior of the equi-separation law observed in multilayer neural networks [16]. At model initialization, the PR of contextualized token embeddings for next-token prediction may exhibit an upward trend from lower to higher layers. However, after a certain amount of training (e.g., around 8,000 steps), the law of equi-learning becomes apparent. Beyond this point, the decay ratio continues to decrease until convergence, with the law consistently manifesting during this phase.
@@ 2 Perspectives from the Law
Fig. 3: Pythia-1B [3] trained on the Pile dataset for various training steps, using a batch size of 2 million, with the total number of steps reaching 143,000. Refer to Fig. S7 in the Supplementary Materials for an enlarged version at initialization (Step=0).
Fig. 4: A 2.8B GPT-2 model pre-trained on a 4 billion token subset of C4 over multiple epochs.
Fig. 5: A 2.8B GPT-2 model pre-trained on varying numbers of unique tokens from the C4 dataset, with a constant total of 55 billion training tokens. Note that the number of unique tokens equals 55 billion divided by the number of repeats.
@@
Fig. 6: Illustration of the law of equi-learning with varying model sizes. Please note that the x axis and y axis share the same ranges within the same model series (same row).
Model scaling.
It is well established that larger models generally yield improved performance [22]. In this part, we explore the impact of model scaling on the observed law. As depicted in Fig. 6, our analysis reveals four consistent trends across three model series, each based on a distinct network architecture: 1) larger models tend to exhibit reduced PR for first-layer token embeddings; 2) larger models tend to show a decrease in PR for last-layer token embeddings; 3) larger models demonstrate an increased layer-wise decay ratio ($\rho$); and 4) larger models exhibit a reduced overall decay ratio, defined as the PR of last-layer token embeddings divided by the PR of first-layer token embeddings. The first trend may be attributed to the increased size or improved quality of first-layer token embeddings. The other trends suggest that while larger models typically display enhanced feature learning capabilities, resulting in more refined last-layer token embeddings, their performance at the individual layer level may be less effective compared to that of smaller models. These findings indicate that the law of equi-learning provides a more nuanced perspective on LLM behavior, particularly in the context of model scaling, offering insights that extend beyond the limitations of test loss alone.
@@
Fig. 7: Different probing tasks are used for BERT (uncased), RoBERTa, and T5. Under the next-token prediction (NTP) task, the law of equi-learning does not appear because these models are pre-trained on different tasks. For BERT and RoBERTa, under their pre-training task, masked language modeling (MLM), the law of equi-learning appears but is somewhat noisy. In contrast, for T5, even with its pre-training task of span corruption (SC), the law of equi-learning does not appear.
Pre-training task.
In addition to the prevalent next-token prediction task, various other pre-training tasks have been employed in the development of LLMs, including masked language modeling [7] and span corruption [33] (for clarity, these three tasks are abbreviated as NTP, MLM, and SC, respectively). As shown in Fig. 7, different probing tasks are used to analyze the contextualized token embeddings of BERT [7], RoBERTa [25], and T5 [33]. Initially, we employ the mainstream NTP task to examine BERT, RoBERTa, and T5. However, the law of equi-learning does not emerge, possibly due to differences in their pre-training tasks compared to NTP. We then apply their respective pre-training tasks—MLM for BERT and RoBERTa, and SC for T5—to probe the models. For BERT and RoBERTa, under MLM, the law of equi-learning appears but is noisy, likely due to the complexity of the token replacement strategy used during pre-training. Specifically, 15% of tokens were masked and replaced with 80% [MASK] tokens, 10% random tokens, and 10% unchanged tokens for the purpose of predicting the original tokens. For T5, even with SC, the law does not emerge, which may be attributed to its encoder-decoder architecture. In SC, the decoder input often lacks natural coherence and relies heavily on the encoder input, while the decoder’s cross-attention layer might further complicate the learning of contextualized token embeddings. These findings suggest that the choice of pre-training task is critical for the emergence of the law of equi-learning, with more naturalistic tasks potentially facilitating its appearance. This may provide supporting evidence for the superiority of NTP, the currently dominant pre-training task.
@@
Fig. 8: The contextualized token embeddings of the current token ($x_{t}$) at each layer are utilized to predict various tokens in the sequence, ranging from previous token ($x_{t-1}$) to next next token ($x_{t+2}$), including the default next token ($x_{t+1}$). It is observed that LLMs tend to forget previous information (positive correlations between the PR and the layer index), including the current token, and improve their prediction of future tokens (negative correlations between the PR and the layer index) after the learning of contextualized token embeddings across layers.
Information flow.
Controlling the flow of information is crucial for effective feature learning in sequential data [17, 4]. To elucidate the dynamics of information flow in LLMs across different architectures, we analyze the contextualized token embeddings of the current token ($x_{t}$) at each layer and their ability to predict other tokens within the sequence, ranging from the previous token ($x_{t-1}$) to the next-next token ($x_{t+2}$), including the default next token ($x_{t+1}$). As depicted in Fig. 8, our results reveal a clear pattern: following the learning of contextualized token embeddings across layers, LLMs exhibit a tendency to forget prior information, including the current token itself, as indicated by the positive correlations between PR and layer index. In contrast, the prediction of future tokens improves, evidenced by negative correlations between PR and layer index. These findings suggest that as the learning of contextualized token embeddings progresses from lower to higher layers, LLMs increasingly discard historical information while simultaneously enhancing their predictive capabilities for upcoming tokens.

### v3

@@ Abstract
Large language models (LLMs) have been widely employed across various application domains, yet their black-box nature poses significant challenges to understanding how these models process input data internally to make predictions. In this paper, we introduce a precise and quantitative law that governs the learning of contextualized token embeddings through intermediate layers in pre-trained LLMs for next-token prediction. Our findings reveal that each layer contributes equally to enhancing prediction accuracy, from the lowest to the highest layer—a universal phenomenon observed across a diverse array of open-source LLMs, irrespective of their architectures or pre-training data. We demonstrate that this law offers new perspectives and actionable insights to inform and guide practices in LLM development and applications, including model scaling, pre-training tasks, and interpretation.
1 Introduction
@@
The primary difficulty arises from the hierarchical nature of LLM architectures, such as Transformer [45], RWKV [31], and Mamba [14]. These architectures are composed of multiple layers, each corresponding to simple functions such as linear or quadratic transformations—better known as the attention mechanism—or a nonlinear combination of both. While the input and output can be observed and specified, the internal transformation of data by each layer becomes elusive due to the hierarchical composition. Specifically, in the case of generative pre-trained transformers (GPT) [32], it is unclear how the embeddings of the input text are progressively transformed into features across different layers for next-token prediction, where a token refers to a word or subword. This generative nature of LLMs introduces additional complexity compared to traditional classification tasks in multilayer perceptrons (MLP): the vocabulary size typically exceeds the embedding dimension, and models undergo relatively few training epochs. The goal of this paper is to shed light on the inner workings of LLMs as they process token embeddings across all layers. In particular, we aim to identify universal and quantitative patterns that can provide useful principles and insights to refine training processes and enhance interpretability in LLMs.
Fig. 1: The law of equi-learning in large language models. Throughout the paper, the x axis represents the layer index, and the y axis, on a logarithmic scale, represents the prediction residual (PR) defined in Eq. 2, unless otherwise specified. The Pearson correlation coefficients, by row first, are -0.997, -0.994, -0.994, -0.988; -0.983, -0.993, -0.992, -0.988; -0.991, -0.989, -0.988, -0.989; -0.988, -0.994, -0.993, -0.992; -0.992, -0.991, -0.991, -0.997. More details can be found in the Supplementary Materials.
Fig. 2: Intermediate-layer contextualized token embeddings for patients</w> (red), cells</w> (blue), and disorder</w> (green) plotted on the plane of the first two principal components. The embeddings are extracted from the GPT-1 model using $200$ sentences sampled from the MedRAG-Textbooks dataset [49] in the medicine domain. As the model progresses from lower to upper layers, the contextualized token embeddings exhibit a clear and progressively increasing separation. The x-axis and y-axis represent the first and second principal components, respectively. More details can be found in the Supplementary Materials.
@@ 2 Main Results
LLMs are nonlinear models designed to predict the subsequent token given a sequence of preceding tokens. Formally, a model processes an input sequence of tokens $x_{1},x_{2},\ldots,x_{t},\ldots$. Each token $x_{t}$ is initially mapped to a vector $\mathbf{h}_{t,0}$ in the embedding space. These embeddings subsequently undergo transformation through a hierarchical stack of model layers—typically comprising attention mechanisms and various operations—yielding a sequence of contextualized token embeddings $\mathbf{h}_{t,\ell}$ at each layer $1\leq\ell\leq L$. Notably, $\mathbf{h}_{t,\ell}$ is computed using the outputs from the preceding layer $\{\mathbf{h}_{j,\ell-1}\mid 1\leq j\leq t\}$, which, due to the autoregressive nature of LLMs, contains information exclusively from tokens at the current position and all preceding positions. Finally, the LLM uses the embedding from the last layer, denoted $\mathbf{h}_{t,\text{last}}:=\mathbf{h}_{t,L}$, to predict the subsequent discrete token $x_{t+1}$. Aggregating all such pairs across the entire training corpus yields the dataset $\mathcal{D}:=\{(\mathbf{h}_{t,\text{last}}^{s},x^{s}_{t+1})\mid 1\leq s\leq S\}$.
@@
where the sum is over all $x_{\text{next}}=x^{s}_{t+1}$, $\hat{x}_{\text{next}}=\mathbf{w}\cdot\mathbf{h}_{t,\text{last}}^{s}+b$, and $\bar{x}_{\text{next}}$ represents the mean of all $x^{s}_{t+1}$. In statistical terms, this measure is known as the fraction of variance unexplained, equivalent to one minus the coefficient of determination [47]. It serves as a canonical metric for evaluating the proportion of variance in the dependent variable that remains unaccounted for by the independent variables. A high PR value indicates limited predictive power of the token embeddings, while a low value suggests strong predictive capability. Thus, PR inherently captures how effectively the token embeddings explain next-token prediction [^1] See more elaboration on this metric in Section 4., aligning with the linear probing paradigm widely employed to analyze the structural properties of contextualized representations in LLMs [18, 26].
To investigate how the predictive power of an LLM with depth $L$ evolves across its layers, we calculate the PR for the next-token prediction task at each intermediate layer. Let $\text{PR}_{l}$ denote this value for the $l$-th layer, where $1\leq l\leq L$. Specifically, instead of using the last-layer embedding $\mathbf{h}_{t,\text{last}}\equiv\mathbf{h}_{t,L}$, we use the embedding of the current token at layer $\ell$, denoted $\mathbf{h}_{t,\ell}$, to predict the next token when computing $\text{PR}_{\ell}$. The law of equi-learning (see Fig. 1) states that the dynamics of predictive power across layers follow the relationship
@@
for some decay ratio $0<\rho<1$. Since the token embeddings at the input layer (the 0-th layer) are not yet contextualized, they are not included. This implies that $\log\text{PR}_{l+1}-\log\text{PR}_{l}\approx-\log\frac{1}{\rho}$, indicating a roughly constant reduction in the logarithm of the PR value across each layer, hence the name equi-learning. The Pearson correlation coefficients between the logarithm of the PR value and the layer index range from 0.983-0.983 to 0.997-0.997 in Fig. 1. For GPT-1, for example, the decay ratio $\rho$ is approximately $0.993$ (as shown in the top-left plot of Fig. 1). In general, $\rho$ depends on various factors, including the model architecture, pre-training data, model depth, feature dimension, and pre-training time.
@@
This law provides what may be the first precise geometric characterization of the learning process for contextualized token embeddings within the intermediate layers of LLMs. Notably, the pre-training objective focuses solely on the last-layer embeddings, aiming to minimize a loss function associated with the last-layer value of PR. Surprisingly, this training dynamics inherently ensure that each layer contributes equally, rather than allowing some layers to carry a disproportionate amount of the workload.
While related phenomena have been observed in MLPs for classification tasks [17, 30], the law presented in this work makes distinct contributions through its focus on LLMs trained for next-token prediction. This task fundamentally differs from classification in several key aspects: it processes sequential data with progressively increasing text lengths and operates within an inherently probabilistic framework without ground truth labels for token predictions. Furthermore, whereas previous studies primarily examined the terminal phase of MLP training, our work demonstrates that the law of equi-learning manifests well before the terminal phase in LLMs, which typically undergo very few epochs of training. This early emergence is particularly noteworthy given the substantial complexity of Transformer-based architectures compared to MLPs. Importantly, this law illuminates the internal mechanisms of the sophisticated LLM architectures [48].
@@
The universality of this law is demonstrated by its emergence across a diverse range of open-source LLMs, spanning different architectures, model sizes, and pre-training datasets. Our investigation employs a comprehensive collection of probing datasets to calculate PR and evaluate the law of equi-learning, including BookCorpus [52], C4 [35], OpenWebText [13], Wikipedia [11], peS2o [39], The Pile [12], Redpajama [7], and OSCAR [40].
Fig. 3: Pythia-1B [3] trained on the Pile dataset for various training steps, using a batch size of 2 million, with the total number of steps reaching 143,000. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2. Refer to Fig. S7 in the Supplementary Materials for an enlarged version at initialization (Step=0).
@@
When does the law emerge? To deepen our understanding of the law’s dynamics during training, we investigate the effects of three key factors—training steps, training epochs, and data repetition—on its emergence throughout the process. As illustrated in Fig. 3, the progression of this law as a function of training steps closely resembles the behavior of the equi-separation law observed in MLPs [17]. At model initialization, the PR of contextualized token embeddings for next-token prediction may exhibit an upward trend from lower to higher layers. However, after a certain amount of training (e.g., around 8,000 steps), the law of equi-learning becomes apparent. Beyond this point, the decay ratio continues to decrease until convergence, with the law consistently manifesting during this phase.
Fig. 4: A 2.8B GPT-2 model pre-trained on a 4 billion token subset of C4 over multiple epochs. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
Fig. 5: A 2.8B GPT-2 model pre-trained on varying numbers of unique tokens from the C4 dataset, with a constant total of 55 billion training tokens. Note that the number of unique tokens equals 55 billion divided by the number of repeats. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
@@ 3 Perspectives from the Law
Fig. 6: Illustration of the law of equi-learning with varying model sizes. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2. Please note that the x axis and y axis share the same ranges within the same model series (same row).
Model scaling. It is well established that larger models generally yield improved performance [24]. In this part, we explore the impact of model scaling on the observed law. As depicted in Fig. 6, our analysis reveals four consistent trends across three model series, each based on a distinct network architecture: (1) larger models exhibit lower PR values for first-layer token embeddings; (2) larger models demonstrate smaller PR values at the final layer, indicating enhanced next-token prediction capabilities; (3) larger models demonstrate an increased layer-wise decay ratio ($\rho$); and (4) larger models exhibit a reduced overall decay ratio, defined as the PR of last-layer token embeddings divided by the PR of first-layer token embeddings. The first trend may be attributed to the increased size or improved quality of first-layer token embeddings. The other trends suggest that while larger models typically display enhanced feature learning capabilities, resulting in more refined last-layer token embeddings, their performance at the individual layer level may be less effective compared to that of smaller models. These findings indicate that the law of equi-learning provides a more nuanced perspective on LLM behavior, particularly in the context of model scaling, offering insights that extend beyond the limitations of test loss alone.
@@
Fig. 7: Different probing tasks are used for BERT (uncased), RoBERTa, and T5. Under the next-token prediction (NTP) task, the law of equi-learning does not appear because these models are pre-trained on different tasks. For BERT and RoBERTa, under their pre-training task, masked language modeling (MLM), the law of equi-learning appears but is somewhat noisy. In contrast, for T5, even with its pre-training task of span corruption (SC), the law of equi-learning does not appear. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2. Note that the prediction target in PR is changed from the next token to the masked token for MLM and the corrupted span for SC.
Pre-training task. In addition to the prevalent next-token prediction task, various other pre-training tasks have been employed in the development of LLMs, including masked language modeling [8] and span corruption [35] (for clarity, these three tasks are abbreviated as NTP, MLM, and SC, respectively). As shown in Fig. 7, different probing tasks are used to analyze the contextualized token embeddings of BERT [8], RoBERTa [27], and T5 [35]. Initially, we employ the mainstream NTP task to examine BERT, RoBERTa, and T5. However, the law of equi-learning does not emerge, possibly due to differences in their pre-training tasks compared to NTP. We then apply their respective pre-training tasks—MLM for BERT and RoBERTa, and SC for T5—to probe the models. For BERT and RoBERTa, under MLM, the law of equi-learning appears but is noisy, likely due to the complexity of the token replacement strategy used during pre-training. Specifically, 15% of tokens were masked and replaced with 80% [MASK] tokens, 10% random tokens, and 10% unchanged tokens for the purpose of predicting the original tokens. For T5, even with SC, the law does not emerge, which may be attributed to its encoder-decoder architecture. In SC, the decoder input often lacks natural coherence and relies heavily on the encoder input, while the decoder’s cross-attention layer might further complicate the learning of contextualized token embeddings. These findings suggest that the choice of pre-training task is critical for the emergence of the law of equi-learning, with more naturalistic tasks potentially facilitating its appearance. This may provide supporting evidence for the superiority of NTP, the currently dominant pre-training task.
@@
Fig. 8: The contextualized token embeddings of the current token ($x_{t}$) at each layer are utilized to predict various tokens in the sequence, ranging from previous token ($x_{t-1}$) to next next token ($x_{t+2}$), including the default next token ($x_{t+1}$). It is observed that LLMs tend to forget previous information (positive correlations between the PR and the layer index), including the current token, and improve their prediction of future tokens (negative correlations between the PR and the layer index) after the learning of contextualized token embeddings across layers. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2. Note that the prediction target in PR is changed from the next token to the previous token, current token, or next next token in settings other than next-token prediction.
Information flow. Controlling the flow of information is crucial for effective feature learning in sequential data [19, 5]. To elucidate the dynamics of information flow in LLMs across different architectures, we analyze the contextualized token embeddings of the current token ($x_{t}$) at each layer and their ability to predict other tokens within the sequence, ranging from the previous token ($x_{t-1}$) to the next-next token ($x_{t+2}$), including the default next token ($x_{t+1}$). Specifically, the embedding of the current token at layer $\ell$, denoted as $\mathbf{h}_{t,\ell}$, is used to predict not only the next token ($x_{t+1}$), but also the previous token ($x_{t-1}$), the current token ($x_{t}$), and the next next token ($x_{t+2}$) when computing $\text{PR}_{\ell}$. As depicted in Fig. 8, our results reveal a clear pattern: following the learning of contextualized token embeddings across layers, LLMs exhibit a tendency to forget prior information, including the current token itself, as indicated by the positive correlations between PR and layer index. In contrast, the prediction of future tokens improves, evidenced by negative correlations between PR and layer index. These findings suggest that as the learning of contextualized token embeddings progresses from lower to higher layers, LLMs increasingly discard historical information while simultaneously enhancing their predictive capabilities for upcoming tokens.